OpenStudy (anonymous):

12(2x - 8) = 4(5 - 2x) + 3(7 - 3x)

OpenStudy (precal):

distribute all of the numbers in front of the parenthesis

OpenStudy (precal):

combine like terms

OpenStudy (tennistar):

apply distribution first 24x - 96 = 20 - 8x + 21 - 9x then combine coefficients of x on the left, and constants on the right 41x = 137 divide by 41 on both sides x = 137/41